OpenStudy (anonymous):

Lets say i know the substitution method, but i cant use it on a test, how do i find the antiderivative of a function that involves the chain rule?

OpenStudy (anonymous):

give an example!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\int\limits_{}^{} ((5x+8)^.5+11)/\sqrt5x+8)\]

OpenStudy (anonymous):

the .5 obviously represent raised to 1/2, and the whole term 5x+8 to the right of the dividion sign is under the square root

OpenStudy (anonymous):

u can use substitution here, substitue 5x+8 power half as something, then its derivative will have 1/root of(5x+8).. then you can easily integrate

OpenStudy (anonymous):

sorry for the late reply

OpenStudy (anonymous):

Yeah, he is saying that the class hasn't done it yet, and he's concerned he will be 'marked down' for using substitution to solve integrals that involve the chain rule. However there really is no other way to solve them, so either you will not be given integrals which involve the chain rule, or you will be allowed to use substitution.

OpenStudy (anonymous):

but cant we divide

OpenStudy (anonymous):

In that case you could do the division yes: \[\int \frac{\sqrt{5x + 8} + 11}{\sqrt{5x+8}}dx = \int 1 + \frac{11}{\sqrt{5x+8}}dx\] But how will you solve the rightmost term without chain rule?

OpenStudy (anonymous):

you can.. lol its direct integral.. sorry i thought you wanted to use substitution!!

OpenStudy (anonymous):

right side is of the the form 1/2rootx whos integral is root x

OpenStudy (anonymous):

well, i think we can factor out the 11, and be left with sqrt5x+8

OpenStudy (anonymous):

then intergrate from there, giving us x+22(5x+8)^1/2

OpenStudy (anonymous):

Nerp.

OpenStudy (anonymous):

the second term will have 5 in the denominator!

OpenStudy (anonymous):

Still need the chain rule because you'll be missing a factor of 5.

OpenStudy (anonymous):

we can divide by 5

OpenStudy (anonymous):

oh, wait i am still getting a 5 in the denominator

OpenStudy (anonymous):

Err blah

OpenStudy (anonymous):

where are you getting a 5 in the denominator if it's not from the chain rule?

OpenStudy (anonymous):

after i integrate i get x+22(5x+8)^1/2

OpenStudy (anonymous):

, so now i want ot check that this is right so i take the derivative

OpenStudy (anonymous):

how should i ajust for the cahin rule

OpenStudy (anonymous):

If 22(5x+8)^1/2 is the integral of 11/sqrt(5x+8) then the derivative of 22(5x+8)^1/2 should be 11/sqrt(5x+8).

OpenStudy (anonymous):

i am not sure, did you figure out the integral

OpenStudy (anonymous):

Yes. But you have to use a substitution.

OpenStudy (anonymous):

no way to do it without

OpenStudy (anonymous):

\[Let\ u = 5x+8\]\[\implies du = 5dx\]\[\implies dx = \frac{1}{5}du\] \[11\int \frac{1}{\sqrt{5x+8}}dx = 11\int\frac{1}{\sqrt{u}}(\frac{1}{5})du\]

OpenStudy (anonymous):

Nope, you cannot avoid it just like you cannot avoid the chain rule when you take the derivative.

OpenStudy (anonymous):

man, i remeber my prof, doin git in classs witout substitution, i got o go look at my notes

OpenStudy (anonymous):

Nope. If you try you will be missing the factor of 1/5.